What is the maximum amount in moles of P2O5 that can theoretically be made from 112 g of O2 and excess phosphorus?
Express your answer to three significant figures and include the appropriate units.
1
Expert's answer
2020-04-15T01:52:38-0400

"P_4+5O_2\to2P_2O_5"

moles of "O_2=\frac{112}{32}mol"

5 moles of "O_2" form 2 moles of "P_2O_5"

so "\frac{112}{32}moles" will form "\frac{2}{5}\times\frac{112}{32}=\frac{7}{5}mol=1.4moles" of "P_2O_5"